# Multi-layer diffusion model of photovoltaic installations

### v1

@@ Abstract
Nowadays, harmful effects of climate change are becoming increasingly apparent. A vital issue that must be addressed is the generation of energy from non-renewable and often polluted sources. For this reason, the development of renewable energy sources is of great importance. Unfortunately, too rapid spread of renewables can disrupt stability of the power system and lead to energy blackouts. One should not simply support it, without ensuring sustainability and understanding of the diffusion process. In this research, we propose a new agent-based model of diffusion of photovoltaic panels. It is an extension of the $q$-voter model that utilizes multi-layer network structure. The model is analyzed by Monte Carlo simulations and mean-field approximation. The impact of parameters and specifications on the basic properties of the model is discussed.
@@
diffusion of innovation , mean-field approximation , Monte Carlo simulations , multi-layer networks
@@ 1 Introduction
Nowadays, we are experiencing effects of climate change and environmental pollution more and more often. What might have once seemed insignificant, such as glaciers melting thousands of kilometers away, now begins to affect us directly. Perhaps the most frightening, as well as tragic, natural disasters of recent years have been the gigantic fires consuming not only warm–climate countries such as Australia or Greece, but also those that used to be cold – Canada [1] or Sweden [2]. These are neither the only cases nor impossible scenarios for other countries. Although many have fortunately not experienced a similar environmental disaster so far, lot of them, including the one of the authors’ origin, Poland, face a different threat almost every day – air pollution. In Poland, apart from industry, low-quality heating installations in single-family households or multi-family tenements are the main causes of its formation. Generally speaking, the problem is the generation of energy from non-renewable and often polluted sources. High emissions of pollutants directly threaten our lives and, in the already coming future, lead to dangerous climate change [3]. According to some studies, global warming may lead to aggravation and spread of many dangerous diseases [4].
@@
That is why the development of renewable energy sources (renewables, RES) is so important. Across the European Union in 2023, about $45\%$ of electricity generation came from RES [5]. Poland, although it saw a rapid expansion of renewables, still falls behind with around $26\%$ [6]. In Europe, wind is the main source of renewable energy [7]. Unfortunately, wind farms require huge investments and vast free space, far from residential buildings. An alternative devoid of these limitations, and rapidly gaining popularity in Poland over the past few years, are photovoltaic panels (PVs). See Fig. 1 for exact numbers. Those can be installed on a roof of a single-family house without adversely affecting the quality of life, while the cost of such a project is achievable already for a middle-income family [8]. However, it is not only important to provide this opportunity and encourage it through subsidy programs or public campaigns, but to ensure sustainability as well. Excessive support of RES diffusion can lead to a significant increase in the variability of demand for conventional generation on a 24-hour basis (the so-called ,,duck curve”), and in extreme cases can even cause loss of stability of the power system. Examples are countries/regions with a large operation of the sun during a year, such as Australia [9] or California [10], but also Poland’s neighbor – Germany [11]. Apart from that, there is another threat, more closely related to the common folk. Present transmission system grids in Poland and other countries are not prepared for such a rapid PVs expansion. If unrestricted, it may and already did lead to grid overload and blackouts [12].
@@
Figure 1: Data on solar installations in Poland in years 2019–2023, according to ARE (Polish energy market agency). More detailed data is being collected from 2021 onwards, most likely due to the growing interest in the prosumer market. Data freely available at www.are.waw.pl.
@@
The aim of this research is to better understand how various factors impact the diffusion of PVs. Diffusion is understood here as the process of spreading a new product through a population. Modeling such a phenomenon can be, and initially was, done in a very simplified manner. A classic example is the Bass diffusion model – consisting of a single differential equation [13]. However, for a problem as complex as the diffusion of PV installations, deterministic models are just not enough [14]. The aforementioned Bass diffusion of innovation model can dramatically change its adoption process, when rewritten to an agent-based model (ABM) and tested on a network structure [15]. In social sciences, an agent-based model is usually understood as a simulation of certain interactions between so-called agents (representing individuals, households, companies, etc.), and taking place within a certain structure, symbolizing a network of acquaintances, contacts or cooperation [16]. Mathematically, we would call it a graph, with vertices being the agents, and edges – the connections between them. ABMs allow for a much more accurate representation of reality, including uniqueness of individuals and the interactions between them [17]. As such, they are currently one of the most powerful tools in studying opinion dynamics and diffusion of innovation [18]. Although they have been used for years to model the diffusion of new energy solutions [19, 20, 21, 22, 23], their applications in modelling the diffusion of PVs remain few [24, 25, 26, 27].
@@
Although much more advanced than simple deterministic models, ABMs are still merely a hypothetical approximation of reality. With that in mind, one should adjust them to a problem at hand as accurately as possible, instead of constructing a one-size-fits-all model and calling it a day. For example, when modeling a spread of gossip in a high school class, a simple network of class acquaintances would be sufficient. However, to properly represent the flow of information and exchange of opinions in the modern world, a much more complex structure is needed. This is where multi-layer networks come in. They are applied in many fields of science [28], for they can provide multi-level representation of real world dependencies [29]. For instance, an individual (agent) may learn about recent sport results either from friends at work (one layer), or through social media (another layer).
@@
Sociologists have long pointed out that structures of social interaction should not be reduced to single-layer networks. However, multi-layer structures have only been studied intensively for a few years [30]. Recently, they have been used in modeling the diffusion of innovation [31, 32]. However, this is a fresh concept. One issue that arises with multi-layer networks is generalization of models’ rules that where originally implemented on single-layer structures. For instance, one may assume that social influence is only effective if it comes from all the layers (AND rule). However, it can also be assumed that the influence is effective even if it comes from a single layer only (OR rule) [33]. In this research, we follow the approach from [34] and study both these variants.
@@
Monte Carlo computer simulation is the main research method for ABMs. However, in some limited cases, as for networks being complete or random graphs with a low clustering coefficient, analytical methods such as Mean-Field (MFA) or Pair Approximation (PA) can be used [35, 36, 37]. Unfortunately, real world networks are characterized by a high clustering coefficient [17], so often the analytical results obtained by these differ quantitatively from the results obtained through Monte Carlo simulations [36]. Although analytical methods are inferior to Monte Carlo simulation in numerous aspects, they possess one clear advantage over the latter – low computational requirements. The need to sweep a multidimensional space of input parameters and perform multiple independent repetitions to obtain good statistical parameters makes Monte Carlo simulations very time-consuming. For this reason, we utilize both computer simulations and analytical methods in this research.
@@
A popular agent-based model that already has found its use in studying diffusion of eco-innovation is the $q$-voter model [19, 22, 38]. In this model, conformity is the basic form of social response. Agents are characterized by a single binary variable denoting their opinion and placed in nodes of some underlying graph structure. Their opinions change upon impact with so-called groups of influence, i.e. $q$ of the agent’s neighbors chosen randomly, but only if a given group presents an unanimous opinion. A well-established extension to the $q$-voter model is the addition of independence – a probability that an agent acts and changes opinion independently of the group of influence [39]. The $q$-voter model has been already examined on multi-layer networks, but to the best of our knowledge only ones consisting of two mirror layers [34, 40].
@@
In this paper, we introduce a new model based on the aforementioned $q$-voter model with independence. Our goal is to capture both opinion dynamics and diffusion of innovation jointly. Hence, we introduce (compared to the original $q$-voter model) a second agent’s attribute, in addition to the opinion – an adoption state. Similarly to the opinion, it is a binary one. Positive value means that an agent possesses a PV installation, negative – it does not. An agent (in this case a household) can acquire knowledge of PVs from two sources. Either it sees panels on the roofs of neighboring households, or communicates with friends/colleagues. To model this twofold dynamic, we utilize a 2-layer network as the underlying topology. The first layer of the network depicts the spacial location of an agent, similarly to [41, 42]. From a mathematical point of view it is just a square lattice (SL) with Moore’s neighborhood [43]. At this level, each agent possesses only knowledge of the adoption states of the closest neighborhood, and these adoption states shape agent’s opinion, according to the $q$-voter rule. The agent cannot share opinion or be subject to the opinions of the neighborhood. This is the visual observation part. A single-family house owner sees only PV panels, or lack thereof, on the roofs of neighboring houses, and does not exchange information or opinions with neighbors. The second layer of the network corresponds to structure of contacts and relationships of agents. For this reason, here we use 2-dimensional Watts-Strogatz graph (WS2D) [44]. It possesses a characteristics of real world social networks [17, 45] and, in a specific case (when randomness equals 00), reduces to a square lattice. At this level, a connection between two agents implies their familiarity, through which they can exchange opinions, but do not observe their adoption states. In this case, social influence is also described by the $q$-voter model. Lastly, it is an agent’s opinion solely that impacts its adoption state. An agent with a positive opinion can install photovoltaic panels with probability $a_{1}$, one with a negative one can resign with probability $a_{2}$. The reason for this research is to thoroughly examine this hypothetical model, to understand how its parameters shape the outcome, and finally, to reproduce the growth of installed power observed in the real data (Fig. 1).

### v2

@@ Abstract
Nowadays, harmful effects of climate change are becoming increasingly apparent. A vital issue that must be addressed is the generation of energy from non-renewable and often polluting sources. For this reason, the development of renewable energy sources is of great importance. Unfortunately, too rapid spread of renewables can disrupt stability of the power system and lead to energy blackouts. One should not simply support it, without ensuring sustainability and understanding of the diffusion process. In this research, we propose a new agent-based model of diffusion of photovoltaic panels. It is an extension of the $q$-voter model that utilizes multi-layer network structure. The novelty is studying both opinion dynamics and diffusion of innovation simultaneously, on a multidimensional structure. The model is analyzed with Monte Carlo simulations and mean-field approximation. The impact of parameters and specifications on the basic properties of the model is discussed. Firstly, we show that for a certain range of parameters, innovation always succeeds, regardless of the initial conditions. Secondly, that the mean-field approximation gives qualitatively the same results as computer simulations, even though it does not utilize knowledge of the network structure.
@@
agent-based modeling , complex networks , computational statistics , diffusion of innovation , dynamical system , opinion dynamics , mathematical modeling , renewable energy
@@ 1 Introduction
Nowadays, we are experiencing effects of climate change and environmental pollution more and more often. What might have once seemed insignificant, such as glaciers melting thousands of kilometers away, now begins to affect us directly. While the gigantic fires that devastated countries like Australia, Canada or Greece have spared the one of the author’s origin, Poland, we experience another environmental disaster – air pollution. In Poland, low-quality heating installations in single-family households or multi-family tenements are among the main causes of its formation [1]. Generally speaking, the problem is the generation of energy from non-renewable and often polluting sources. High emissions of pollutants directly threaten our lives and, in the already coming future, lead to dangerous climate change [2], of which global warming is probably the most widely discussed. It not only affects entire ecosystems and the comfort of life, but according to some studies may lead to aggravation and spread of many dangerous diseases as well [3].
@@
That is why the development of renewable energy sources (renewables, RES) is so important. Across the European Union in 2023, about $45\%$ of electricity generation came from RES [4]. Poland, although it saw a rapid expansion of renewables, still falls behind with around $26\%$ [5]. In Europe, wind is the main source of renewable energy [6]. Unfortunately, wind farms require huge investments and vast free space, far from residential buildings. An alternative devoid of these limitations, and rapidly gaining popularity in Poland over the past few years, are photovoltaic panels (PVs), see Fig. 1 for exact numbers. Those can be installed on a roof of a single-family house without adversely affecting the quality of life, while the cost of such a project is achievable already for a middle-income family [7]. However, it is not only important to provide this opportunity and encourage it through subsidy programs or public campaigns, but to ensure sustainability as well. Excessive support of RES diffusion can lead to a significant increase in the variability of demand for conventional generation on a 24-hour basis, and in extreme cases can even cause loss of stability of the power system. Examples are countries/regions with a large operation of the sun during a year, such as Australia [8] or California [9], but also Poland’s neighbor – Germany [10]. Apart from that, there is another threat, more closely related to the common folk. Present transmission system grids in Poland and other countries are not prepared for such a rapid PVs expansion. If unrestricted, it may lead to grid overload and blackouts. In fact, this is an already emerging risk [11].
@@
Figure 1: Data on solar installations in Poland in years 2019–2023, according to ARE (Polish energy market agency). More detailed data is being collected from 2021 onward, most likely due to the growing interest in the prosumer market. Data freely available at www.are.waw.pl.
@@
The aim of this research is to better understand how various factors impact the diffusion of PVs. Diffusion is understood here as the process of spreading a new product through a population [12]. Diffusion of innovation has been modeled for years. It began in a very simplified manner. A classic example is the Bass diffusion model – a fully deterministic one, consisting of a single differential equation [13]. However, for a problem as complex as the diffusion of PV installations, aggregated models are just not enough [14]. They are unable to describe clustering of individuals, a phenomenon observed in real life [15]. The aforementioned Bass diffusion of innovation model can dramatically change its behavior, when rewritten to an agent-based model (ABM) and tested on a network structure [16]. In social sciences, an agent-based model is usually understood as a simulation of certain interactions between so-called agents (representing individuals, households, companies, etc.), and taking place within a certain structure, symbolizing a network of acquaintances, contacts or cooperation [17]. Mathematically, we would call such a structure a graph, with vertices being the agents, and edges – the connections between them [18]. ABMs allow for a much more accurate representation of reality, including heterogeneity of individuals and the interactions between them [18]. As such, ABMs are currently one of the most powerful tools in studying opinion dynamics and diffusion of innovation [19]. Although they have been used for years to model the diffusion of new energy solutions [20, 21, 22, 23, 24], their applications in modeling the diffusion of PVs remain few [25, 26, 27, 28].
@@
Although much more advanced than simple deterministic models, ABMs are still merely a hypothetical approximation of reality. With that in mind, one should adjust them to a problem at hand as accurately as possible, instead of constructing a one-size-fits-all model. A huge role here is played by the underlying structure. For example, when modeling a spread of gossip in a high school class, a simple network of class acquaintances would be sufficient. However, to properly represent the flow of information and exchange of opinions in the modern world, a much more complex structure is needed. This is where multi-layer networks come in. They are applied in many fields of science [29], for they can provide multi-level representation of real world dependencies [30]. For instance, an individual (agent) may learn about recent sport results either from friends at work (one layer), or through social media (another layer).
@@
Sociologists have long pointed out that structures of social interaction should not be reduced to single-layer networks [30]. However, multi-layer structures have only been studied intensively since the last decade [31]. Recently, they have been used in modeling the diffusion of innovation [32, 33]. Nevertheless, this is still a relatively fresh concept. One issue that arises with multi-layer networks is generalization of models’ rules that where originally implemented on single-layer structures. For instance, one may assume that social influence is only effective if it comes from all the layers (AND rule). However, it can also be assumed that the influence is effective even if it comes from a single layer only (OR rule) [34]. In this research, we follow the approach from [35] and study both variants.
@@
Monte Carlo (MC) computer simulations are the main research method for ABMs. However, in some limited cases, as for networks being complete or random graphs with a low clustering coefficient, analytical methods such as mean-field (MFA) or pair approximation (PA) can be used [36, 37, 38]. Unfortunately, real world networks are characterized by a high clustering coefficient [18], so often the analytical results obtained this way differ quantitatively from ones given by Monte Carlo simulations [37]. The need to sweep a multidimensional space of input parameters and perform multiple independent repetitions to obtain good statistical parameters makes Monte Carlo simulations very time-consuming. For this reason, we utilize both computer simulations and analytical methods in this research.
@@
A popular agent-based model that already has found its use in studying diffusion of eco-innovation is the $q$-voter model [20, 23, 39]. In this model, conformity is the basic form of social response. Agents are characterized by a single binary variable denoting their opinion and placed in nodes of some underlying graph structure. Their opinions change upon impact with so-called groups of influence, i.e. $q$ of the agent’s neighbors chosen randomly, but only if a given group presents an unanimous opinion. A well-established extension to the $q$-voter model is the addition of independence – a probability that an agent acts and changes opinion independently of the group of influence [40]. The $q$-voter model has been already examined on multi-layer networks, but to the best of our knowledge only ones consisting of two identical layers [35, 41].
@@
In this paper, we introduce a new model based on the aforementioned $q$-voter model with independence. Our goal is to capture both opinion dynamics and diffusion of innovation jointly. Hence, we introduce (compared to the original $q$-voter model) a second agent’s attribute, in addition to the opinion – an adoption state. Similarly to the opinion, it is a binary one. Positive value means that an agent possesses a PV installation, negative – it does not. An agent (in this case a household) can acquire knowledge of PVs from two sources. Either it sees panels on the roofs of neighboring households, or communicates with friends/colleagues. To model this twofold dynamic, we utilize a two-layer network as the underlying topology. The first layer of the network depicts the spacial location of an agent, similarly to [42, 43]. From a mathematical point of view it is just a square lattice (SL) with Moore’s neighborhood [44]. At this level, each agent possesses only knowledge of the adoption states of the closest neighborhood, and these adoption states shape agent’s opinion, according to the $q$-voter rule. The agent cannot share opinion or be subject to the opinions of the neighborhood. This is the visual observation part. A single-family house owner sees only PV panels, or lack thereof, on the roofs of neighboring houses, and does not exchange information or opinions with neighbors. The second layer of the network corresponds to structure of contacts and relationships of agents. For this reason, here we use two-dimensional Watts-Strogatz graph (WS2D) [45]. It possesses some characteristics of real world social networks [18, 46] and, in a specific case (when randomness equals 00), reduces to a square lattice. At this level, a connection between two agents implies their familiarity, through which they can exchange opinions, but do not observe their adoption states. In this case, social influence is also described by the $q$-voter model. Lastly, it is an agent’s opinion solely that impacts its adoption state. An agent with a positive opinion can install photovoltaic panels with probability $a_{1}$, one with a negative one can resign with probability $a_{2}$. The reason for this research is to thoroughly examine this mathematical model, to understand how its parameters shape the outcome.